# MANet: Fine-Tuning Segment Anything Model for Multimodal Remote Sensing Semantic Segmentation

### v1

@@ IV-F Model Scale Analysis
The improved performance of MANet is largely attributable to the general knowledge provided by the vision foundation model, SAM, which is also a large model. However, the large model does not have advantages in terms of computational complexity or inference speed compared to existing general methods. Consequently, we focus on reporting the model’s trainable parameter number and memory footprint to measure its hardware requirements.
@@
Table VI presents the results of the model scale for all methods compared in this work. As indicated in Table VI, the proposed parameter-efficient fine-tuning technique allows the large foundation models to be used on a single GPU while maintaining a manageable number of trainable parameters and memory costs. In our experiments, we successfully fine-tuned the ViT-L backbone on the same hardware with the same hyperparameters, achieving results that surpassed all existing methods. For the ViT-H backbone, we adjusted the batch size from $10$ to $6$ due to the GPU memory limitation. This reduction in batch size did not degrade performance, but further improved performance. These results prove the powerful feature extraction and fusion capabilities of the large vision foundation model. This work also offers valuable insights for exploring multimodal tasks with large models under constrained hardware conditions.
TABLE VI: Model scale analysis measured by a $256\times 256$ image on a single NVIDIA GeForce RTX 3090 GPU. MIoU values are the results on the Vaihingen dataset. The best results are in Bold.
Method Parameter (M) Memory (MB) MIoU (%)
ABCNet [43] 13.39 1598 75.20
PSPNet [47] 46.72 3124 76.96
MAResU-Net [44] 26.27 1908 79.89
UNetFormer [29] 24.20 1980 81.97
RS3Mamba [50] 43.32 1548 82.78
TransUNet [49] 93.23 3028 78.26
FuseNet [9] 42.08 2284 78.71
vFuseNet [30] 44.17 2618 78.92
ESANet [45] 34.03 1914 79.42
SA-GATE [48] 110.85 3174 81.27
CMFNet [14] 123.63 4058 81.44
MFTransUNet [15] 43.77 1549 81.61
CMGFNet [46] 64.20 2463 82.26
FTransUNet [16] 160.88 3463 84.23
MANet (ViT-B) 20.43 2274 82.71
MANet (ViT-L) 56.67 5514 84.36
MANet (ViT-H) 111.28 8934 85.03
@@ V Conclusion
In this study, a novel multimodal fusion framework called MANet has been proposed for semantic segmentation of multimodal remote sensing data by leveraging the general knowledge embedded in the vision foundation model, SAM. Specifically, the SAM’s image encoder is equipped with a novel MMAdapter that is utilized to extract and fuse multimodal remote sensing features. The fused deep features are then further exploited by a pyramid-based DFM before being reconstructed into segmentation maps. Comprehensive experiments on two multimodal datasets, ISPRS Vaihingen and ISPRS Potsdam, have demonstrated that MANet outperforms current state-of-the-art segmentation methods. This research is the first to confirm the reliability of SAM on DSM data and offers a promising solution for applying vision foundation models to multimodal data. Furthermore, it has the potential for extension to other remote sensing tasks, such as semi-supervised or unsupervised learning.

### v2

@@ IV-F Model Scale Analysis
The improved performance of MFNet is largely attributable to the general knowledge provided by the vision foundation model, SAM. However, SAM is also a large model, and the large model does not have advantages in terms of computational complexity or inference speed compared to existing general methods. Consequently, we focus on reporting the model’s trainable parameter number and memory footprint to measure its hardware requirements.
Table VI presents the results of the model scale for all methods compared in this work. As indicated in Table VI, the proposed multimodal fine-tuning techniques allow the large foundation models to be used on a single GPU while maintaining a manageable number of trainable parameters and memory costs. The parameter statistics of MFNet are divided into two parts: the fine-tuning parameters in SAM’s image encoder + the parameters in DFM and the decoder. The parameters in the latter remain consistent across different MFNet configurations. A comparison between MMLoRA and MMAdapter shows that MMLoRA significantly reduces the number of parameters by compressing thousands of dimensional spaces into a rank of $4$ through low-rank decomposition. While this approach is generally efficient, it may result in the loss of some essential information, especially when processing complex remote sensing data. Consequently, MMAdapter outperforms MMLoRA in terms of performance.
@@
In our experiments, we successfully fine-tuned the ViT-L backbone on the same hardware with the same hyperparameters, achieving results that surpassed all existing methods. For the ViT-H backbone, we adjusted the batch size from $10$ to $4$ due to the GPU memory limitation. This reduction in batch size did not degrade performance, but further improved performance. These results prove the powerful feature extraction and fusion capabilities of the large vision foundation model. This work also offers valuable insights for exploring multimodal tasks with large models under constrained hardware conditions.
IV-G Discussion
This work introduces a unified multimodal fine-tuning framework with two SAM-based multimodal fine-tuning mechanisms. As an early exploration of this field, we thoroughly investigate the performance of the vision foundation model on remote sensing multimodal tasks by developing two classical fine-tuning approaches: Adapter and LoRA. Comprehensive analytical experiments are conducted to evaluate these methods. Additionally, MFNet offers a straightforward multimodal fusion network, paving the way for future research in the following directions:
  • Improving fine-tuning modules: This work employs two representative fine-tuning techniques, namely Lora and Adapter, to demonstrate the framework’s effectiveness. Future research is encouraged to apply more advanced variants, such as [lei2023conditional, chen2023hadamard, liu2024dora, hayou2024lora+] to diverse multimodal remote sensing tasks. In particular, more efficient fine-tuning strategies for large-scale models are worth exploring, as foundation models typically require substantial memory resources.
  • Improving fusion modules: The work employs an adaptive weight-based method for feature fusion during the encoding stage. Future work could explore more advanced and effective fusion strategies tailored to MMAdapter and MMLoRA. Similarly, the fusion of deep high-level features can be enhanced with other fusion mechanisms, such as cross-attention, to improve performance.
  • Addressing challenging categories: SAM’s performance on challenging categories, such as distinguishing between very similar trees and low vegetation, or accurately detecting small objects such as cars, demands further investigation. To improve accuracy, it may be necessary to develop specialized object recognition modules that focus on these specific tasks. This could involve category-specific feature extraction techniques.
  • Exploration of other remote sensing modalities: This study demonstrates the superiority of multimodal fine-tuning framework using optical images and DSM data as examples, and also provides valuable insights into the potential of combining these modalities. However, the performance of SAM on other remote sensing modalities, such as multispectral, LiDAR, and SAR, offers an exciting avenue for future investigation. Exploring these modalities could further enhance SAM’s capabilities and broaden its applicability in diverse remote sensing tasks.
Overall, this work serves as a foundational framework, with several aspects left open for further exploration. We hope it can be broadly extended to various types of multimodal remote sensing tasks.
@@ V Conclusion
In this study, we proposed a unified multimodal fusion framework with multimodal fine-tuning for remote sensing semantic segmentation, leveraging the general knowledge embedded in the vision foundation model, SAM. Using two representative single-modal fine-tuning mechanisms, namely Adapter and LoRA, we demonstrated the seamless integration of existing mechanisms into the proposed unified framework for extracting and fusing multimodal features from remote sensing data. The fused deep features were further refined using a pyramid-based DFM and reconstructed into segmentation maps. Comprehensive experiments on three benchmark multimodal datasets, ISPRS Vaihingen, ISPRS Potsdam and MMHunan, confirmed that MFNet achieves superior performance compared to current state-of-the-art segmentation methods. This research represents the first validation of SAM’s reliability on DSM data and offers a promising pathway for leveraging vision foundation models in multimodal remote sensing tasks. Moreover, the proposed framework has the potential to be extended to other remote sensing applications, including semi-supervised and unsupervised learning tasks.